OpenStudy (anonymous):

∫ (5 / x² -4/x) dx I'm confused as to what to do. Do I begin by taking the anti derivative and if so how it is done?

OpenStudy (anonymous):

5x^-2 - 4x^-1 took the anti derivative and got 5x^-1/-1 - 4x^0/0. Do I ignore the second term since it is undefined?

OpenStudy (kropot72):

The second term in your putative answer is \[\infty\]. Then you should add the constant of integration as the third term of the answer.

OpenStudy (anonymous):

Im confused as to how an integer divided by zero is equivilant to infinity?

OpenStudy (kropot72):

0/0 is indeterminate (undefined). However any real number divided by zero will give the result of infinity - subject to a positive or negative sign.

OpenStudy (kropot72):

Sorry. The second term in your putative answer should be: \[-(4\times \ln x)\]

OpenStudy (anonymous):

ohh okay that makes much more sense

OpenStudy (callisto):

∫ (5 / x² -4/x) dx = ∫ (5x^-2 -4/x) dx =∫ (5x^-2 )dx -∫ (4/x) dx = 5 [x^(-2+1)] / (-2+1) -4ln|x| +C = -5 /x - 4ln|x| +C

OpenStudy (kropot72):

You're welcome. The integral of 1/x does not follow the usual rule for integration.

OpenStudy (anonymous):

Thanks guys I really appreciate youre help!